# TypedThinker: Typed Thinking Improves Large Language Model Reasoning

### v1

@@ 5 Conclusion and Limitation
In this paper, we investigate how reasoning types diversify LLMs’ thinking and propose a novel framework TypedThinker to incorporate different reasoning types into problem-solving. TypedThinker is inspired by human cognition processes during reasoning: it learns an implicit policy to select the appropriate reasoning types with the meta-thinker and to conduct the selected type of reasoning with the reasoner. It also maintains an explicit memory to retrieve experiences to aid reasoning. TypedThinker optimizes its implicit policy and the explicit memory with its own successful experiences during the self-training process. The results show that TypedThinker enhances the reasoning capabilities of Mistral 7B and LLaMA3 8B on four benchmarks. Furthermore, TypedThinker shows good generalization capabilities in new domains and models. It can also improve GPT-4o’s performance by the effective reasoning type selection.
Despite the promising results, TypedThinker has several limitations that need further investigation. Firstly, one problem may require different reasoning types at different steps, and applying one sole reasoning type can hardly find a correct solution. In that case, dividing the problems into multiple reasoning steps, and applying TypedThinker for each step could make the reasoning more diverse and effective. Additionally, this paper mainly focuses on logical and mathematical benchmarks. Expanding to a broader range of tasks, such as commonsense reasoning or creative problem-solving, could deepen the understanding of the role of reasoning types in various problems and provide a more comprehensive assessment of TypedThinker’s capabilities.
@@ A.2 Dataset Processing
For GSM8k and MATH, we follow  Toshniwal et al. (2024) to process the dataset.
A.3 Analysis of Reasoning Types
Accuracy of Reasoning Types We calculate the accuracy for each reasoning type on our empirical dataset $\overline{D}$, shown in Figure 6. We can find that on LogiQA and MATH, the accuracy of different reasoning types is similar. However, deductive and analogical reasoning outperform the other two on BBH while inductive and abductive reasoning are more effective. The results illustrate that after our carefully designed demonstration for each reasoning type, LLM’s capabilities in other reasoning types achieve comparable performance with deductive reasoning. This ensures the quality and the balance of our collected dataset on each reasoning type.
@@ A.3 Analysis of Reasoning Types
Diversity of Reasoning Types To further verify whether the reasoning types can make the solutions more diverse, we compare the diversity between solutions under different sampling settings in Table 10. We use Levenshtein Distance (Levenshtein, 1966) and the n-gram overlaps between sentences to evaluate diversity. Specifically, for $K$ generations $G=\{g_{1},\cdots,g_{K}\}$ of the same problem, we calculate the distance between each pair and normalize them with the sentence length. Then the average distance over these paired results is used as the distance of these $K$ generations. If we denote the normalized Levenshtein Distance function as $f_{\rm{ld}}$, this process can be represented as:
@@
LogiQA Zero-shot @ 5 0.3043 0.5877 0.5367
Zero-shot + 5 types 0.5997 0.2584 0.1863
Few-shot @ 5 0.5729 0.2316 0.1226
Few-shot + 5 types 0.6437 0.1745 0.0773
BBH Zero-shot @ 5 0.5170 0.3104 0.2157
Zero-shot + 5 types 0.7117 0.1280 0.0495
Few-shot @ 5 0.5992 0.2239 0.1191
Few-shot + 5 types 0.6495 0.1750 0.0764
GSM8k Zero-shot @ 5 0.6242 0.1951 0.0907
Zero-shot + 5 types 0.6831 0.1513 0.0540
Few-shot @ 5 0.4977 0.3117 0.1761
Few-shot + 5 types 0.7097 0.1366 0.0476
MATH Zero-shot 0.6726 0.1573 0.0703
Zero-shot + 5 types 0.7291 0.1116 0.0352
Few-shot 0.6588 0.1741 0.0804
Few-shot + 5 types 0.7319 0.1154 0.0382
A.4 Impact Statement
@@
Abduction John is 24 years younger than his dad. The sum of their ages is 68 years. How many years old is John? Retrieval: Question: Lisa is 10 years younger than her mom. The sum of their ages is 70 years. How old is Lisa?
@@
Analogy The integer m is between 30 and 80 and is a multiple of 6. When m is divided by 8, the remainder is 2. Similarly, when m is divided by 5, the remainder is 2. What is the value of m? To solve this problem using abductive reasoning, we assume that one possible value of m exists that abides by the constraints and check if this assumption holds.1. Fi rst, filter values of m that are multiples of 6 between 30 and 80. 2. Next, apply the condition that when m is divided by 8, the remainder is 2. Only 42, 66 fit this condition.3. Apply the third condition, that when divided by 5, m should leave a remainder of 2. Testing the applicable values so far and find 42 meets the requirement. So the answer is 424242.

### v2

@@ 5 Conclusion and Limitation
We investigate how reasoning types diversify LLMs’ thinking and propose TypedThinker to incorporate different reasoning types into problem-solving. TypedThinker is inspired by human cognition processes during reasoning: it learns an implicit policy to select the appropriate reasoning types with the meta-thinker and to apply the selected type of reasoning with the reasoner. It also maintains an explicit memory to retrieve experiences to aid reasoning. The results show that TypedThinker enhances the reasoning capabilities of Mistral 7B, LLaMA3 8B and Qwen 2 7B on four benchmarks. Furthermore, TypedThinker shows good generalization capabilities in new domains and models.
Despite the promising results, TypedThinker has several limitations that need further investigation. Firstly, one problem may require different reasoning types at different steps, and applying one sole reasoning type can hardly find a correct solution. In that case, dividing the problems into multiple reasoning steps, and applying TypedThinker for each step could make the reasoning more diverse and effective. Additionally, this paper mainly focuses on logical and mathematical benchmarks. Expanding to a broader range of tasks, such as code generation and creative problem-solving, could deepen the understanding of the role of reasoning types in various problems and provide a more comprehensive assessment of TypedThinker’s capabilities.
@@ A.2.1 Data Processing
GSM8k (Cobbe et al., 2021) is a commonly used math benchmark to evaluate LLMs’ capability in math reasoning. It contains 8.5K grade school math word problems, which are split into 7.5k training examples and 1k test problems. Each problem usually takes between 2 and 8 steps to solve. MATH (Hendrycks et al., 2021) is also a popular math benchmark for LLMs. It contains 12,500 challenging competition mathematics problems with 7 categories. There are 7.5k training examples and 5k test problems. We follow  Toshniwal et al. (2024) to process the dataset.
Contexthub (Hua et al., 2024) is a new propositional logic benchmark. It contains abstract and contextualized logical problems from 12 categories with 4 levels of difficulty (Zhu et al., 2024). We follow the standard split of the original paper and use the subset of difficult level 4 to test the complex logic reasoning capabilities. The abstract logical problems only contain the symbolic variable without natural language description, which can be viewed as symbolic reasoning problems.
Livebench (White et al., 2024) is a recently proposed benchmark with 18 diverse tasks across 6 categories, specifically designed to minimize data contamination. All problems have verifiable, objective ground-truth answers, allowing hard questions to be scored accurately and automatically. We evaluate our models on three tasks (spatial, web of lies v2, zebra puzzle) from the reasoning category, splitting them 0.7/0.3 for training and testing.
A.2.2 Dataset Examples
We demonstrate one example for each dataset below.
LogiQA One seminar had 18 participants. It is known that :(1) At least 5 young teachers are female; (2) At least 6 female teachers are over middle age; (3) At least 7 young women are teachers; According to the above information, which can be concluded? Options: (A) Some young teachers are not women (B) Some young women are not teachers (C) There are at least 11 young teachers (D) There are at least 13 female teachers
BBH: logical deduction three objects The following paragraphs each describe a set of three objects arranged in a fixed order. The statements are logically consistent within each paragraph. In a golf tournament, there were three golfers: Ada, Mel, and Mya. Mya finished below Ada. Mel finished above Ada. Options: (A) Ada finished last (B) Mel finished last (C) Mya finished last
GSM8k A 40 meters rope was cut into 2 parts in the ratio of 2:3. How long is the shorter part?
MATH: Algebra Level 1 $361+2(19)(6)+36=x$. Solve for $x$.
ContextHub: Abstract - Level 2 (wqeq or mnze) → zkx. (NOT ttjmx) → kottz. (kottz or zkx) → pofk. Given pofk is False, what is the value of ttjmx?
LiveBench: reasoning - zebra puzzle There are 3 people standing in a line numbered 1 through 3 in a left-to-right order. Each person has a set of attributes: Nationality, Music-Genre, Transport. The attributes have the following possible values: - Nationality: spanish, argentine, canadian - Music-Genre: punk, rock, reggae - Transport: train, jet-ski, trike and exactly one person in the line has a given value for an attribute. Given the following premises about the line of people: - the person who is argentine avoids getting on a train - the person who is spanish is somewhere between the person who listens to punk and the person who listens to rock - the person who listens to punk is not anywhere to the right of the person that travels by trike - the person who listens to punk is on the immediate right of the person that travels by jet-ski Answer the following question: What is the nationality of the person who listens to rock? Return your answer as a single word, in the following format: ***X***, where X is the answer.
A.3 Training Details
A.4 Analysis of Typed Reasoning
Accuracy of Typed Reasoning We calculate the accuracy for each reasoning type on our empirical dataset $\overline{D}$, shown in Figure 6. We can find that on LogiQA and MATH, the accuracy of different reasoning types is similar. However, deductive and analogical reasoning outperform the other two on BBH while inductive and abductive reasoning are more effective. The results illustrate that after our carefully designed demonstration for each reasoning type, LLM’s capabilities in other reasoning types achieve comparable performance with deductive reasoning. This ensures the quality and the balance of our collected dataset on each reasoning type.
@@ A.4 Analysis of Typed Reasoning
Diversity of Typed Reasoning To further verify whether the reasoning types can make the solutions more diverse, we compare the diversity between solutions under different sampling settings in Table 10. We use Levenshtein Distance (Levenshtein, 1966) and the n-gram overlaps between sentences to evaluate diversity. Specifically, for $K$ generations $G=\{g_{1},\cdots,g_{K}\}$ of the same problem, we calculate the distance between each pair and normalize them with the sentence length. Then the average distance over these paired results is used as the distance of these $K$ generations. If we denote the normalized Levenshtein Distance function as $f_{\rm{ld}}$, this process can be represented as:
@@
LogiQA Zero-shot @ 5 0.304 0.588 0.537
Zero-shot + 5 types 0.600 0.258 0.186
Few-shot @ 5 0.573 0.232 0.123
Few-shot + 5 types 0.644 0.175 0.077
BBH Zero-shot @ 5 0.517 0.310 0.216
Zero-shot + 5 types 0.712 0.128 0.050
Few-shot @ 5 0.599 0.224 0.119
Few-shot + 5 types 0.650 0.175 0.076
GSM8k Zero-shot @ 5 0.624 0.195 0.091
Zero-shot + 5 types 0.683 0.151 0.054
Few-shot @ 5 0.498 0.312 0.176
Few-shot + 5 types 0.710 0.137 0.048
MATH Zero-shot 0.673 0.157 0.070
Zero-shot + 5 types 0.729 0.112 0.035
Few-shot 0.659 0.174 0.080
Few-shot + 5 types 0.732 0.115 0.038
A.5 More experimental results
A.5.1 Results on More Backbone LLMs
We conducted further experiments using Qwen 2-7B-Instruct (Bai et al., 2023) as our backbone LLM. The Qwen series of open-source large language models have demonstrated comparable or even superior performance to the Mistral and LLaMA families across multiple tasks. The results are shown in Table 11. Our method achieves approximately 7% improvement over the few-shot baseline in both single-generation and majority-vote settings (+SC @5). These results demonstrate that TypedThinker is a general and effective method for enhancing the reasoning capabilities of various LLMs.
Table 11: Qwen 2-7B-Instruct results. The annotations are the same with Table 1.
LogiQA BBH GSM8K MATH Avg.
Few-shot 0.552 0.471 0.646 0.417 0.521
+ SC @ 5 0.579 0.554 0.763 0.497 0.598
CoT Selection 0.554 0.516 0.772 0.451 0.573
+ SC @ 5 0.560 0.528 0.780 0.497 0.591
Zeroshot MoR 0.573 0.498 0.492 0.407 0.493
Fewshot MoR 0.589 0.581 0.889 0.551 0.652
TypedThinker 0.595 0.534 0.779 0.474 0.596
+ SC @ 5 0.644 0.584 0.880 0.565 0.668
A.5.2 Results on more benchmarks
We further evaluate our methods on LiveBench (White et al., 2024) to test the generalization capability of our method. The experimental settings are the same as described in Section 4.5. The results are shown in Table 12. It demonstrates that TypedThinker outperforms other baselines, further supporting its generalization capability across diverse tasks.
Table 12: TypedThinker outperforms other baselines on LiveBench without extra finetuning. Here the results are based on the majority vote over 5 responses (+SC @5).
Mistral 7B LLaMA3 8B
Few-shot 0.178 0.200
CoT Selection 0.244 0.200
TypedThinker 0.267 0.267
A.5.3 More Ablation Studies
The primary reason for comparing our method with the few-shot baseline is that fine-tuning for specific reasoning types is an integral part of our approach. Therefore, we evaluate the impact of our fine-tuned reasoner through a separate ablation study. However, comparing TypedThinker to few-shot baselines without fine-tuning may not fully account for the benefits of fine-tuning. Therefore, we conduct two more experiments to verify the influence of the fine-tuned LLMs.
Comparison with base LLM + one module Ablation studies in Section 4.4 investigate the contribution of each component by removing one component each time. Here we provide additional ablation results by adding one component to the base LLM each time, resulting in two variants: Base LLM + Meta-thinker and Base LLM + Collection. For a more reliable conclusion, we ran experiments three times to calculate the average and std and present the result in Table 13 and 14.
Table 13: Mistral 7B results are based on three repetitive experiments. Avg. indicates the average accuracy over four benchmarks.
LogiQA BBH GSM8K MATH Avg.
Few-shot 0.493 $\pm$ 0.007 0.347 $\pm$ 0.01 0.372 $\pm$ 0.014 0.071 $\pm$ 0.003 0.321 $\pm$ 0.006
CoT Selection 0.475 $\pm$ 0.009 0.361 $\pm$ 0.01 0.377 $\pm$ 0.011 0.104 $\pm$ 0.008 0.329 $\pm$ 0.004
LLM + Meta Thinker 0.512 $\pm$ 0.003 0.377 $\pm$ 0.006 0.379 $\pm$ 0.004 0.106 $\pm$ 0.009 0.343 $\pm$ 0.004
LLM + Collection 0.519 $\pm$ 0.007 0.398 $\pm$ 0.005 0.363 $\pm$ 0.004 0.086 $\pm$ 0.008 0.342 $\pm$ 0.001
TypedThinker 0.553 $\pm$ 0.004 0.430 $\pm$ 0.008 0.390 $\pm$ 0.012 0.103 $\pm$ 0.01 0.369 $\pm$ 0.006
Table 14: LLaMA 3 8B results are based on three repetitive experiments. Avg. indicates the average accuracy over four benchmarks.
LogiQA BBH GSM8K MATH Avg.
Few-shot 0.569 $\pm$ 0.003 0.319 $\pm$ 0.006 0.476 $\pm$ 0.004 0.102 $\pm$ 0.005 0.366 $\pm$ 0.001
CoT Selection 0.558 $\pm$ 0.011 0.376 $\pm$ 0.007 0.360 $\pm$ 0.006 0.104 $\pm$ 0.009 0.349 $\pm$ 0.007
LLM + Meta Thinker 0.538 $\pm$ 0.005 0.434 $\pm$ 0.005 0.508 $\pm$ 0.005 0.118 $\pm$ 0.005 0.400 $\pm$ 0.001
LLM + Collection 0.574 $\pm$ 0.011 0.497 $\pm$ 0.004 0.438 $\pm$ 0.005 0.109 $\pm$ 0.006 0.404 $\pm$ 0.001
TypedThinker 0.546 $\pm$ 0.004 0.534 $\pm$ 0.003 0.535 $\pm$ 0.001 0.203 $\pm$ 0.009 0.455 $\pm$ 0.002
Results show that the retrieval component improves performance on logical tasks but may mislead models on mathematical datasets. This is consistent with our findings in the ablation study in Table 2: the retrieved solutions with digits may mislead the model. Meanwhile, compared with the ICL reasoner, our finetuned reasoner shows better capability in identifying the suitable reasoning type.
A.6 Discussion on More Reasoning Problems
In this paper, we mainly focus on logical and math reasoning problems. However, our TypedThinker can also be extended to symbolic or commonsense reasoning without extra efforts. For example, the problems under abstract category in ContextHub can be viewed as symbolic reasoning, as shown in A.2.2. Therefore, in addition to the overall performance shown in Table 4, we present specific performance on the abstract category of symbolic reasoning in Table 15. As we can see TypedThinker outperforms baseline methods, even without further fine-tuning the meta-thinker and reasoner for this symbolic reasoning task.
Table 15: TypedThinker performs best on the abstract category of Contexthub. Here the results are based on the majority vote over 5 responses (+SC @5).
Mistral 7B LLaMA3 8B
Few-shot 0.25 0.35
CoT Selection 0.55 0.40
Zero-shot MoR 0.55 0.35
Few-shot MoR 0.45 0.45
Self-Discover 0.25 0.55
TypedThinker 0.75 0.55
While the LogiQA dataset contains problems requiring commonsense reasoning, the dataset lacks explicit annotations (such as a specific category) for such tasks. For example, here is one case that requires commonsense knowledge about manufacturing costs, market dynamics, and consumer preferences.
LogiQA: A commonsense reasoning example Traditionally, the most highly sought cars have been the sports cars and similar two-door models. Nevertheless, Zincstone Motors has chosen to eliminate the last two-door models and produce only four-door models. Which of the following would, if true, most help to explain Zincstone Motors’ strategy? Options: (A) In almost every instance, Zincstone Motors models lead all comparable models of competitors in fuel efficiency and have lower average maintenance costs as well. (B) After a spate of recent additional safety requirements, the cost of frame and doors of Zincstone Motors’ standard two-door models are now three times as expensive as standard four-door frame and doors. (C) Many of Zincstone Motors models are exported and sold overseas, including in some countries like Japan, which import a significant number of cars into the United States. (D) As American consumers lose access to car manufacturers who produce two-door cars, and as two-door cars occupy smaller and smaller shares of the United States car market, American consumers’ tastes tend to shift from two-door cars.
For reasoning problems that significantly differ from the existing domains (logic and math), additional demonstrations tailored to the task are recommended to guide reasoning. For example, inductive reasoning might involve deriving a general program from exemplar input-output test cases in code generation (Shao et al., 2024; Yang et al., 2024). In such cases, providing a few task-specific examples or conducting light fine-tuning on the reasoner can further enhance performance.
A.7 Impact Statement
@@
Abduction The integer m is between 30 and 80 and is a multiple of 6. When m is divided by 8, the remainder is 2. Similarly, when m is divided by 5, the remainder is 2. What is the value of m? To solve this problem using abductive reasoning, we assume that one possible value of m exists that abides by the constraints and check if this assumption holds.1. Fi rst, filter values of m that are multiples of 6 between 30 and 80. 2. Next, apply the condition that when m is divided by 8, the remainder is 2. Only 42, 66 fit this condition.3. Apply the third condition, that when divided by 5, m should leave a remainder of 2. Testing the applicable values so far and find 42 meets the requirement. So the answer is 424242.
Analogy John is 24 years younger than his dad. The sum of their ages is 68 years. How many years old is John? Retrieval: Question: Lisa is 10 years younger than her mom. The sum of their ages is 70 years. How old is Lisa?